# The Assouad dimension of Kakeya sets in $\mathbb{R}^{3}$

### v1

@@ Abstract
The above results follow from certain multi-scale structure theorems for arrangements of tubes and rectangular prisms in three dimensions, and a mild generalization of the sticky Kakeya theorem previously proved by the authors.
@@ 1 Introduction
where $\overline{\mathrm{dim}\,}_{M}$ denotes the upper Minkowski dimension, and the infimum is taken over all decompositions $E=\bigcup_{i}E_{i}$ into countably many sets. We have $\mathrm{dim}\,_{P}(E)\leq\mathrm{dim}\,_{H}(E)$ for every set $E\subset\mathbb{R}^{n}$, and if $\mathrm{dim}\,_{P}(E)=\mathrm{dim}\,_{H}(E)=\alpha$, then for every $\varepsilon>0$ there exists a constant $C_{\varepsilon}$ and a measure $\mu$ supported on $E$ so that for $\mu$-a.e. $x\in E$ we have the Frostman-type estimate
$C_{\varepsilon}^{-1}r^{\alpha+\varepsilon}\leq\mu(B(x,r))\leq C_{\varepsilon}r^{\alpha-\varepsilon}\quad\textrm{for all}\ 0<r\leq 1.$
@@
We say a set compact set $K\subset\mathbb{R}^{n}$ is a Kakeya set with stably equal Hausdorff and packing dimension if for all $\varepsilon>0$, there is a measure $\mu$ supported on $K$ with the following properties.
  • There is a constant $C_{\varepsilon}$ so that for all $x\in\operatorname{supp}(\mu),$ we have
    $C_{\varepsilon}^{-1}r^{\alpha+\varepsilon}\leq\mu(B(x,r))\leq C_{\varepsilon}r^{\alpha-\varepsilon},\quad\textrm{for all}\ 0\leq r\leq 1,$
@@
Theorems 1.2 and 1.4 will follow from a common discretized Kakeya estimate, which is the main technical contribution of this paper. In the statement below, $N_{\rho}(X)$ denotes the $\rho$-neighborhood of the set $X$, a $\delta$-tube is the $\delta$-neighborhood of a unit line segment, and say two $\delta$-tubes are essentially distinct if neither tube is contained in the 2-fold dilate of the other; see Section 2 for precise definitions.
@@
In the above theorem, we have required that the tubes be essentially distinct and that every convex set $U\subset\mathbb{R}^{3}$ contain at most $\delta^{-\eta}|U|(\#\mathbb{T})$ tubes from $\mathbb{T}$. This condition is satisfied, for example, if #$\#\mathbb{T}=\delta^{-2}$ and the tubes in $\mathbb{T}$ point in $\delta$-separated directions. Indeed; for the latter condition we need only consider convex sets $U\subset B(0,1)\subset\mathbb{R}^{3}$ of diameter roughly 1, and every set of this form is comparable to an ellipsoid whose axes have lengths $s\leq t\leq w$ with $w\sim 1$. If $T$ is a $\delta$-tube contained in $U$, then the direction of $T$ lies in a rectangular sector of $S^{2}$ of dimensions roughly $s\times t$; if the tubes in $\mathbb{T}$ point in $\delta$-separated directions, we conclude that #$\#\{T\in\mathbb{T}\colon T\subset U\}\lesssim st\delta^{-2}\lesssim|U|(\#\mathbb{T})$. The anti-clustering condition imposed in Theorem 1.5 and will be discussed further in Section 2.2.
@@ 1.1 Main ideas, and a sketch of the proof
Recall that every Kakeya set contains a two-dimensional family of unit line segments pointing in different directions. In the discretized setting, this corresponds to a set of $\delta$-tubes pointing in $\delta$-separated directions. The requirement that the line segments (or $\delta$-tubes) point in different directions can be thought of as a type of anti-clustering condition—without this condition, for example, we could construct a two-dimensional family of coplanar line segments in $\mathbb{R}^{3}$, and the union of these line segments would have dimension at most 2.
In [17], Wolff imposed an anti-clustering condition on his collection of $\delta$-tubes in $\mathbb{R}^{3}$ that is slightly weaker than requiring the tubes to point in $\delta$-separated directions. First, Wolff required that for $\delta\leq\rho\leq 1$, each $\rho$-tube must contain at most $(\rho/\delta)^{2}$ $\delta$-tubes; if we think of $\delta$-tubes as points in the Grassmannian of (affine) lines in $\mathbb{R}^{3}$, then Wolff’s first requirement corresponds to a 2-dimensional Frostman non-concentration condition on the set of tubes. Second, Wolff required that for each $\delta\leq\rho\leq 1$, at most $\rho/\delta$ $\delta$-tubes can be contained in any rectangular prism of dimensions $\delta\times\rho\times 1$. This second condition prevents the tubes from clustering into the $\delta$-neighborhood of a plane. These two conditions are now referred to as the Wolff axioms.
In our arguments below, we will consider several anti-clustering conditions that are related to the Wolff axioms. The first one already appeared in the statement of Theorem 1.5; it is a mild generalization of the Wolff axioms, which we call the Convex Wolff Axioms. Next, we say a set of $\delta$-tubes $\mathbb{T}$ satisfies the Convex Wolff Axioms at every scale if for every $\delta\leq\rho\leq 1$, the tubes in $\mathbb{T}$ can be covered by $\rho$-tubes, so that for each $\rho$-tube $T_{\rho}$ in this cover, the portion of $\mathbb{T}$ contained inside $T_{\rho}$ can be re-scaled to create a collection of $(\delta/\rho)$-tubes that satisfy the convex Wolff axioms. Finally, we say a set of tubes $\mathbb{T}$ satisfies a Frostman Condition at exponent $\alpha$ if every $\rho$-tube contains at most $\rho^{\alpha}(\#\mathbb{T})$ tubes from $\mathbb{T}$. Precise versions of these definitions are given in Section 2.2.
@@
In Section 4, we show that if a set of tubes $\mathbb{T}$ satisfies the Convex Wolff Axioms and also satisfies a Frostman condition, then either it must satisfy the Convex Wolff Axioms at every scale, or else the conclusion of Theorem 1.5 must hold for $\mathbb{T}$ (or both). In broad strokes, the argument is as follows. Suppose that $\mathbb{T}$ satisfies a Frostman condition but fails to satisfy the Convex Wolff Axioms at every scale. Then the tubes in $\mathbb{T}$ must cluster into rectangular prisms of dimensions $s\times t\times 1$ for some $\delta\leq s<\!\!<t\leq 1.$ Each such prism is essentially the intersection of the $s$-neighborhood of a plane with the $t$-neighborhood of a line. First, we consider the case that when two such prisms intersect, the corresponding planes make angle $\theta>\!\!>s/t$. Then the union of the $s$-neighborhood of these planes will contain a ball of radius $r=\theta t,$ which is much larger than $s$, and hence (1.1) holds. Second, we consider the case that when two such prisms intersect, the corresponding planes make angle $\leq s/t$. This forces the $s\times t\times 1$ prisms cluster into rectangular prisms of dimensions $s^{\prime}\times t^{\prime}\times 1$, where $s^{\prime}\ll t^{\prime}$ are substantially larger than $s$ and $t$, respectively. We can then repeat this argument at a larger scale. After finitely many iterations of this argument, we must find ourselves in the First Case. This concludes the argument.
The argument described above is a key geometric input for the proof of Theorem 1.5. With this geometric input, the proof of Theorem 1.5 proceeds by contradiction: suppose there exists a counter-example to Theorem 1.5, i.e. a set of tubes satisfying the Convex Wolff Axioms, for which (1.1) fails for every pair of scales $\rho,r$ and every ball $B$ of radius $r$. The first step is to find a “worst possible” such counter-example to Theorem 1.5, i.e. a set of $\delta$-tubes $\mathbb{T}$ of largest possible cardinality satisfying the Convex Wolff Axioms, for which the reverse of Inequality (1.1) holds for every pair of scales $\rho,r$ and every ball $B$ of radius $r$, with the term $(\rho/r)^{\varepsilon}$ replaced by $(\rho/r)^{\omega},$ for the largest possible exponent $\omega$.
Since the conclusion of Theorem 1.5 is agnostic to location and scale, if we select $s\in[\delta,1]$; consider only those tubes from $\mathbb{T}$ that are contained in a $s$-tube; and re-scale, then the reverse of Inequality (1.1) must continue to hold, with the same exponent $\omega$. But this means that the newly obtained collection of tubes (obtained by restricting $\mathbb{T}$ to a $s$-tube and rescaling) cannot be larger (relative to the new scale $\delta/s$) than the original set $\mathbb{T}$, since $\mathbb{T}$ was a set of largest possible cardinality for which the reverse of Inequality (1.1) holds at every location and scale with exponent $\omega$. The above argument shows that $\mathbb{T}$ must satisfy a Frostman condition. We can now apply the geometric input discussed above (and proved in Section 4): $\mathbb{T}$ satisfies the Convex Wolff Axioms (by assumption), and we just established that $\mathbb{T}$ satisfies a Frostman condition. On the other hand, we are assuming that $\mathbb{T}$ does not satisfy the conclusion of Theorem 1.5. We conclude that $\mathbb{T}$ satisfies the Convex Wolff Axioms at every scale.
In summary, we have shown that if Theorem 1.5 is false, then there must exist a “worst possible” counter-example, and this counter-example must satisfy the Convex Wolff Axioms at every scale. Finally, we arrive at a contradiction (and thus complete the proof of Theorem 1.5) using a slight variant of the Sticky Kakeya theorem from [16]. We state an informal version of this result below.
Theorem 1.8.
For all $\varepsilon>0$, there exists $\eta,\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $\mathbb{T}$ be a set of essentially distinct $\delta$-tubes in $\mathbb{R}^{3}$ that satisfy the Convex Wolff Axioms at every scale. For each $T\in\mathbb{T}$, let $Y(T)\subset T$ be a measurable set, and suppose that $\sum_{\mathbb{T}}|Y(T)|\geq\delta^{\eta}\sum_{\mathbb{T}}|T|$. Then
Theorem 1.8 is closely related to previous work by the authors from [16]. In [16], the authors prove a variant of Inequality (1.3), in which the tubes $\mathbb{T}$ satisfy slightly different hypotheses—the tubes are required to point in different directions and satisfy a Frostman condition at exponent 2. There are two places in [16] where one uses the hypothesis that the tubes point in different directions, and this hypothesis can be replaced by the Convex Wolff Axioms with minimal modification. A precise version of Theorem 1.8 is stated in Section 5. In Section 6 we briefly outline the proof of Theorem 1.8, and how it may be obtained by modifying the arguments in [16].
@@ 1.2 Thanks
The authors would like to thank Jonathan Fraser and Terence Tao for comments and suggestions on an earlier draft of this manuscript. Hong Wang is supported by NSF CAREER DMS-2238818 and NSF DMS-2055544. Joshua Zahl is supported by a NSERC Discovery Grant.
@@ 2 Notation and Definitions
We write $A\lesssim B$ or $A=O(B)$ to mean there is a constant $C$ (which might depend on the ambient dimension $n$) so that $A\leq CB$. We write $A\sim B$ if $A\lesssim B$ and $B\lesssim A$. If the constant $C$ is allowed to depend on an additional parameter such as $\varepsilon$, we then we denote this by $A\lesssim_{\varepsilon}B$. Many of our results will involve a small positive parameter, which we will call $\delta$. We write $A\lessapprox B$ to mean that there is a constant $C$ so that $A\leq C(\log 1/\delta)^{C}B$. As above, if the constant $C$ is allowed to depend on an additional parameter such as $\varepsilon$, we then we denote this by $A\lessapprox_{\varepsilon}B$. Finally, during proof sketches and other informal remarks, we will sometimes write $A\ll B$ to suggest to the reader that that the real number $A$ is much smaller than the real number $B$.
@@ 2.1 Tubes and their shadings
Recall that a $\delta$-tube in $\mathbb{R}^{n}$ is the $\delta$-neighborhood of a unit line segment. We will use $T$ to denote a $\delta$-tube, and $\ell(T)$ to denote its coaxial line. If $T$ is a $\delta$-tube (or more generally, the translation of a centrally symmetric convex set) and $R>0$, we define $RT$ to be the $R$-fold dilate of $T$. We say two tubes $T,T^{\prime}$ are essentially distinct if $T\not\subset 2T^{\prime}$ and $T^{\prime}\not\subset 2T$.
@@
Let $\mathcal{S}$ be a collection of subsets of $\mathbb{R}^{n}$, and let $\tau\in(0,1]$. A collection of sets $\{Y(S)\subset S\colon S\in\mathcal{S}\}$ is called a $\tau$-dense shading of $\mathcal{S}$ if $\sum_{S\in\mathcal{S}}|Y(S)|\geq\tau\sum_{S\in\mathcal{S}}|S|$. $\{Y(S)\}$ is called a uniformly $\tau$-dense shadings of $\mathcal{S}$ if $|Y(S)|\geq\tau|S|$ for each $S\in\mathcal{S}$.
@@
Let $S\subset\mathbb{R}^{n}$ and let $Y(S)$ be a shading. Then there is a regular shading $Y^{\prime}(S)\subset Y(S)$ with $|Y^{\prime}(S)|\geq\frac{1}{2}|Y(S)|$.
@@ 2.2 Wolff axioms and their generalizations
In this section we will define several anti-clustering conditions that we can impose on our collection of tubes. In [17] (see also Definition 13.1 from [7]), Wolff introduced what are now called the Wolff axioms; this is a non-concentration condition that forbids a collection of $\delta$-tubes from clustering near a fatter $\rho$-tube, and also forbids a collection of $\delta$-tubes from concentrating on (affine) 2-planes. The precise definition is as follows.
@@
  • Every $\rho$-tube can contain at most $C(\rho/\delta)^{n-1}$ tubes from $\mathbb{T}$.
  • Every rectangular prism of dimensions $2\delta\times\rho\times 2$ can contain at most $C(\rho/\delta)$ tubes from $\mathbb{T}$.
@@
3. Every set of $\delta$-tubes contained in $B(0,1)\subset\mathbb{R}^{n}$ that satisfies the Wolff axioms has cardinality at most $2C\delta^{1-n}$.
@@
In practice we will have $n=3$, and the sets in $\mathcal{S}$ will be either $\delta$-tubes, or $s\times t\times 1$ rectangular prisms (not necessarily axis-parallel) for some $s,t\in[\delta,1]$.
@@
In our arguments below, we will often consider a Kakeya set (or more accurately a collection of $\delta$-tubes) at many different scales. Typically, for a collection of $\delta$-tubes $\mathbb{T}$, we will want to examine the coarsening of $\mathbb{T}$ at a larger scale $\rho$, and also examine the collection of $\delta$-tubes from $\mathbb{T}$ that are contained in inside each of these coarser $\rho$-tubes. As observed in Remark 2.7 above, if $\mathbb{T}$ satisfies the Convex Wolff Axioms, then the coarsening of $\mathbb{T}$ at a larger scale will also satisfy these axioms [^3] In contrast, the more traditional condition that the tubes in $\mathbb{T}$ point in $\delta$-separated directions is not preserved after coarsening. However, the (rescaled) collection of $\delta$-tubes from $\mathbb{T}$ contained inside each coarser $\rho$-tube might not satisfy the Convex Wolff Axioms. Our next task is to introduce a slightly stricter variant of the Convex Wolff Axioms that addresses this issue.
@@
Let $\mathcal{A}$ be a set and $\mathcal{B}$ be a multi-set of convex subsets of $\mathbb{R}^{n}$. We say that $\mathcal{A}$ is a cover of $\mathcal{B}$ if each set $B\in\mathcal{B}$ is contained in at least one set from $\mathcal{A}$. For each $A\in\mathcal{A}$, we write $\mathcal{B}[A]$ to denote the multi-set $\{B\in\mathcal{B}\colon B\subset A\}$. We say that $\mathcal{A}$ is a $K$-uniform cover of $\mathcal{B}$ if #$\#\mathcal{B}[A]\leq K(\#\mathcal{B}[A^{\prime}])$ for each pair of sets $A,A^{\prime}\in\mathcal{A}$. Finally, we say that $\mathcal{A}$ is a partitioning cover of $\mathcal{B}$ if $\mathcal{A}$ is a cover of $\mathcal{B}$, and for every pair of distinct $A,A^{\prime}\in\mathcal{A}$, the sets $\mathcal{B}[2A]$ and $\mathcal{B}[2A^{\prime}]$ are disjoint, where $2A$ denotes the 2-fold dilate of $A$. In practice, $\mathcal{A}$ and $\mathcal{B}$ will be collections of $\rho$-tubes and $\delta$-tubes respectively, for some $\rho>\delta$.
@@
Let $C\geq 1,\delta>0$. We say a set $\mathbb{T}$ of $\delta$-tubes in $\mathbb{R}^{n}$ satisfies the Convex Wolff Axioms at every scale with error $C$ if the tubes in $\mathbb{T}$ are essentially distinct, and for every $\rho_{0}\in[\delta,1]$, there exists $\rho\in[\rho_{0},C\rho_{0})$ and a set of $\rho$-tubes $\mathbb{T}_{\rho}$ that satisfies the following properties.
@@
    For each $T_{\rho}\in\mathbb{T}_{\rho}$, the unit re-scaling of $\mathbb{T}[T_{\rho}]$ relative to $T_{\rho}$ satisfies the Convex Wolff Axioms with error $C$.
As the name suggests, Definition 2.12 is a multi-scale property. In particular, if $\mathbb{T}$ satisfies the Convex Wolff Axioms at every scale, then we can cover $\mathbb{T}$ by fatter $\rho$-tubes; both the $\rho$-tubes and the (re-scaled) tubes from $\mathbb{T}$ inside each $\rho$-tube will again satisfy the Convex Wolff Axioms at every scale. The precise statement is as follows.

### v2

@@ Abstract
The above results follow from certain multi-scale structure theorems for arrangements of tubes and rectangular prisms in three dimensions, and a generalization of the sticky Kakeya theorem previously proved by the authors.
@@ 1 Introduction
2. Our proof of Theorem 1.2 establishes a slightly stronger (though more technical) result, which asserts that the analogue of Theorem 1.2 continues to hold if we restrict to pairs of scales $\rho<r$ in Definition 1.1 that are “exponentially separated,” in the sense that $\rho<r^{1+\varepsilon}$. This shows that every Kakeya set in $\mathbb{R}^{3}$ has quasi-Assouad dimension 3. (The Assouad dimension of a set is always bounded below by its quasi-Assouad dimension; see [5] for further details).
@@
where $\overline{\mathrm{dim}\,}_{M}$ denotes the upper Minkowski dimension, and the infimum is taken over all decompositions $E=\bigcup_{i}E_{i}$ into countably many sets. We have $\mathrm{dim}\,_{\!P}(E)\geq\mathrm{dim}\,_{\!H}(E)$ for every set $E\subset\mathbb{R}^{n}$, and if $\mathrm{dim}\,_{\!P}(E)=\mathrm{dim}\,_{\!H}(E)=\alpha$, then for every $\varepsilon>0$ there exists a constant $C$ and a measure $\mu$ supported on $E$ so that for $\mu$-a.e. $x\in E$ we have the Frostman-type estimate
$C^{-1}r^{\alpha+\varepsilon}\leq\mu(B(x,r))\leq Cr^{\alpha-\varepsilon}\quad\textrm{for all}\ 0<r\leq 1.$
@@
We say a set compact set $K\subset\mathbb{R}^{n}$ is a Kakeya set with stably equal Hausdorff and packing dimension if for all $\varepsilon>0$, there is a constant $C$ and a measure $\mu$ supported on $K$ with the following properties:
  • For all $x\in\operatorname{supp}(\mu)$ and all $0<r\leq 1$, we have
    $C^{-1}r^{\alpha+\varepsilon}\leq\mu(B(x,r))\leq Cr^{\alpha-\varepsilon},$
@@
Theorems 1.2 and 1.4 will follow from a common discretized Kakeya estimate, which is the main technical contribution of this paper. In the statement below, $N_{\rho}(X)$ denotes the $\rho$-neighborhood of the set $X$; a $\delta$-tube is the $\delta$-neighborhood of a unit line segment; and we say two $\delta$-tubes are essentially distinct if neither tube is contained in the 2-fold dilate of the other. See Section 2 for precise definitions.
@@
In the above theorem, we have required that the tubes be essentially distinct and that every convex set $U\subset\mathbb{R}^{3}$ contain at most $\delta^{-\eta}|U|(\#\mathbb{T})$ tubes from $\mathbb{T}$. This condition is satisfied, for example, if #$\#\mathbb{T}=\delta^{-2}$ and the tubes in $\mathbb{T}$ point in $\delta$-separated directions. Indeed; for the latter condition we need only consider convex sets $U\subset B(0,1)\subset\mathbb{R}^{3}$ of diameter roughly 1, and every set of this form is comparable to an ellipsoid whose axes have lengths $s\leq t\leq w$ with $w\sim 1$. If $T$ is a $\delta$-tube contained in $U$, then the direction of $T$ lies in a rectangular sector of $S^{2}$ of dimensions roughly $s\times t$; if the tubes in $\mathbb{T}$ point in $\delta$-separated directions, we conclude that #$\#\{T\in\mathbb{T}\colon T\subset U\}\lesssim st\delta^{-2}\lesssim|U|(\#\mathbb{T})$. The anti-clustering condition imposed in Theorem 1.5 will be discussed further in Section 2.2.
@@ 1.1 Main ideas, and a sketch of the proof
In [16], the authors proved that every sticky Kakeya set in $\mathbb{R}^{3}$ has Hausdorff and Minkowski dimension 3. In brief, a Sticky Kakeya Set in $\mathbb{R}^{n}$ is a Kakeya set that, when discretized at scale $\delta$, gives rise to a collection of $\delta$-tubes with the following two properties:
  • (a)
    The tubes point in $\delta$-separated directions.
  • (b)
    For each $\delta\leq\rho\leq 1$, the tubes can be covered by a set of $\rho$-tubes, with few (i.e. about 1) $\rho$-tubes pointing in each direction.
Sticky Kakeya sets have two useful properties. First, for $\delta\leq\rho\leq 1$, the $\rho$-neighborhood of a sticky Kakeya set at scale $\delta$ is a sticky Kakeya set at scale $\rho$. In other words, if we thicken a sticky Kakeya set, we obtain a sticky Kakeya set at a coarser scale. Second, for $\delta\leq\rho\leq 1$, the $\rho$-neighborhood of a typical tube in our sticky Kakeya set will contain about $(\rho/\delta)^{1-n}$ $\delta$-tubes; if we rescale this $\rho$-tube to be comparable to the unit ball, then we obtain a sticky Kakeya set at scale $\delta/\rho$. In other words, if we zoom-in and rescale a sticky Kakeya set, we obtain a sticky Kakeya set at a coarser scale.
The Sticky Kakeya Theorem (or more precisely, a generalization that we will prove in Section 6) will play an important role in the proof of Theorem 1.5. In Section 4 we show that every set of $\delta$-tubes satisfying the hypotheses of Theorem 1.5 must either satisfy a suitable analogue of stickiness (and hence we can apply our generalization of the Sticky Kakeya Theorem and obtain the conclusion of Theorem 1.5), or else at least one of the following three things must happen:
  • (i)
    After thickening the tubes at a suitable scale, we obtain a richer collection of tubes, i.e. a collection of tubes that has similar structural properties, but larger cardinality (relative to the new scale).
  • (ii)
    After zooming-in and rescaling at a suitably chosen location and scale, we obtain a richer collection of tubes.
  • (iii)
    After possibly coarsening and zooming-in, the tubes arrange themselves into “planks” (like planks of wood, these sets are long (i.e. length 1) in one direction, much shorter in a second direction, and shorter still in the final direction).
@@
In Section 4 we show that if Item (iii) occurs, then the corresponding collection of tubes satisfies the conclusion of Theorem 1.5. This observation is the main new geometric input in the proof of Theorem 1.5. In brief, the argument is as follows. Suppose each plank has dimensions roughly $s\times t\times 1$, with $\delta\leq s<\!\!<t\leq 1$. This means that each plank has an associated plane, which is the span of the two vectors pointing in the longest and second-longest directions. First we consider the case in which a typical pair of intersecting planks have corresponding planes that make angle $\theta>\!\!>s/t$. If this occurs, then the union of the $s$-neighborhood of these planks will contain a ball of radius $r=\theta t.$ This is much larger than $s$, and hence (1.1) holds. Second, we consider the case in which a typical pair of intersecting planks have corresponding planes that make angle $\leq s/t$. This forces the planks to cluster into bigger planks dimensions $s^{\prime}\times t^{\prime}\times 1$, where $s^{\prime}\ll t^{\prime}\leq 1$ are substantially larger than $s$ and $t$, respectively. We can then repeat this argument at a larger scale. After finitely many iterations of this argument, we must find ourselves in the First Case. This concludes the argument.
Summarizing the arguments above, if $\mathbb{T}$ is a counter-example to Theorem 1.5, then there exists either a thickening or a zooming-in and rescaling that produces a richer collection of tubes. However, Theorem 1.5 is agnostic to location and scale, and this means that if $\mathbb{T}$ is a counter-example to Theorem 1.5, then any thickening or zooming-in and rescaling of $\mathbb{T}$ remains a counter-example. Thus to prove Theorem 1.5, we show that if a counter-example to Theorem 1.5 exists, then there must exist a (nearly) maximally rich counter-example. But this is impossible, since some thickening or zooming-in and rescaling of this (supposedly) maximally rich counter-example is an even richer counter-example.
In Section 2.2 we precisely define the relevant generalization of stickiness, and in Section 3 we define what it means for $\mathbb{T}$ to be a maximally rich counter-example to Theorem 1.5. In Section 4 we carry out the bulk of the geometric arguments described above. In Section 5 we show that a (hypothetical) counterexample to Theorem 1.5 must satisfy the hypotheses of (a generalization of) the sticky Kakeya Theorem. In Sections 6 and 7 we prove the necessary generalization of the Sticky Kakeya Theorem (this is Theorem 5.2), thereby proving Theorem 1.5. Finally, in Section 8 we show how Theorem 1.5 implies Theorems 1.2 and 1.4.
@@ 1.2 Thanks
The authors would like to thank Larry Guth, Jonathan Fraser, Terence Tao, and the anonymous referee for comments, suggestions, and corrections on an earlier draft of this manuscript. The authors would like to thank Sam Craig for pointing out an error in the proof of Lemma 4.4 in an earlier draft of this manuscript. Hong Wang is supported by NSF CAREER DMS-2238818 and NSF DMS-2055544. Joshua Zahl is supported by an NSERC Discovery Grant and an NSERC Alliance Grant.
@@ 2 Notation and definitions
We write $A\lesssim B$ or $A=O(B)$ to mean there is a constant $C$ (which might depend on the ambient dimension $n$) so that $A\leq CB$. We write $A\sim B$ if $A\lesssim B$ and $B\lesssim A$. If the constant $C$ is allowed to depend on an additional parameter such as $\varepsilon$, then we denote this by $A\lesssim_{\varepsilon}B$. Many of our results will involve a small positive parameter, which we will call $\delta$. We write $A\lessapprox B$ to mean that there is a constant $C$ so that $A\leq C(\log 1/\delta)^{C}B$. As above, if the constant $C$ is allowed to depend on an additional parameter such as $\varepsilon$, then we denote this by $A\lessapprox_{\varepsilon}B$. Finally, during proof sketches and other informal remarks, we will sometimes write $A\ll B$ to suggest to the reader that that the real number $A$ is much smaller than the real number $B$.
@@ 2.1 Tubes and their shadings
Recall that a $\delta$-tube in $\mathbb{R}^{n}$ is the $\delta$-neighborhood of a unit line segment. We will use $T$ to denote a $\delta$-tube, and $\ell(T)$ to denote its coaxial line. If $T$ is a $\delta$-tube (or more generally, the translation of a centrally symmetric convex set) and $R>0$, we define $RT$ to be the $R$-fold dilate of $T$ around its center. We say two tubes $T,T^{\prime}$ are essentially distinct if $T\not\subset 2T^{\prime}$ and $T^{\prime}\not\subset 2T$.
@@
Let $\mathcal{S}$ be a collection of subsets of $\mathbb{R}^{n}$, and let $\tau\in(0,1]$. A collection of sets $\{Y(S)\subset S\colon S\in\mathcal{S}\}$ is called a $\tau$-dense shading of $\mathcal{S}$ if $\sum_{S\in\mathcal{S}}|Y(S)|\geq\tau\sum_{S\in\mathcal{S}}|S|$. $\{Y(S),\ S\in\mathcal{S}\}$ is called a uniformly $\tau$-dense shading of $\mathcal{S}$ if $|Y(S)|\geq\tau|S|$ for each $S\in\mathcal{S}$.
@@
Let $\delta>0$, let $S\subset B(0,1)\subset\mathbb{R}^{n}$, and let $Y(S)$ be a shading. Then there is a shading $Y^{\prime}(S)\subset Y(S)$ that is regular at scales $\geq\delta$ and satisfies $|Y^{\prime}(S)|\geq\frac{1}{2}|Y(S)|$.
@@ 2.2 Wolff axioms and their generalizations
In this section we will define several anti-clustering conditions that we can impose on our collection of tubes. In [17] (see also Definition 13.1 from [7]), Wolff introduced what are now called the Wolff axioms; this is a non-concentration condition that forbids a collection of $\delta$-tubes from clustering inside a fatter $\rho$-tube, and also forbids a collection of $\delta$-tubes from clustering near (affine) 2-planes. The precise definition is as follows.
@@
  • Every $\rho$-tube in $\mathbb{R}^{n}$ can contain at most $C(\rho/\delta)^{n-1}$ tubes from $\mathbb{T}$.
  • Every rectangular prism in $\mathbb{R}^{n}$ of dimensions $2\delta\times 2\delta\times\cdots\times 2\delta\times\rho\times 2$ can contain at most $C(\rho/\delta)$ tubes from $\mathbb{T}$.
@@
3. Every set of $\delta$-tubes contained in $B(0,1)\subset\mathbb{R}^{n}$ that satisfies the Wolff axioms has cardinality at most $100^{n}C\delta^{1-n}$.
@@
Informally, when $C$ has size close to 1, we will sometimes suppress the role of $C$ and say that $\mathcal{S}$ satisfies the Convex Wolff Axioms. In practice we will have $n=3$, and the sets in $\mathcal{S}$ will be either $\delta$-tubes, or $s\times t\times 1$ rectangular prisms (not necessarily axis-parallel) for some $\delta\leq s\leq t\leq 1$.
@@
In our arguments below, we will often consider a Kakeya set (or more accurately a collection of $\delta$-tubes) at many different scales. Typically, for a collection of $\delta$-tubes $\mathbb{T}$, we will want to examine the coarsening of $\mathbb{T}$ at a larger scale $\rho$, and also examine the collection of $\delta$-tubes from $\mathbb{T}$ that are contained inside each of these coarser $\rho$-tubes. As observed in Remark 2.7 above, if $\mathbb{T}$ satisfies the Convex Wolff Axioms, then the coarsening of $\mathbb{T}$ at a larger scale will also satisfy these axioms [^3] In contrast, the more traditional condition that the tubes in $\mathbb{T}$ point in $\delta$-separated directions is not preserved after coarsening.. However, the (rescaled) collection of $\delta$-tubes from $\mathbb{T}$ contained inside each coarser $\rho$-tube might not satisfy the Convex Wolff Axioms. Our next task is to introduce a slightly stricter variant of the Convex Wolff Axioms that addresses this issue.
@@
Let $\mathcal{A}$ be a set and $\mathcal{B}$ be a multi-set of convex subsets of $\mathbb{R}^{n}$. We say that $\mathcal{A}$ is a cover of $\mathcal{B}$ if each set $B\in\mathcal{B}$ is contained in at least one set from $\mathcal{A}$. For each $A\in\mathcal{A}$, we write $\mathcal{B}[A]$ to denote the multi-set $\{B\in\mathcal{B}\colon B\subset A\}$. We say that $\mathcal{A}$ is a $K$-uniform cover of $\mathcal{B}$ if #$\#\mathcal{B}[A]\leq K(\#\mathcal{B}[A^{\prime}])$ for each pair of sets $A,A^{\prime}\in\mathcal{A}$. When $K=1$, we will abbreviate this to a uniform cover. Finally, we say that $\mathcal{A}$ is a partitioning cover of $\mathcal{B}$ if $\mathcal{A}$ is a cover of $\mathcal{B}$, and for every pair of distinct $A,A^{\prime}\in\mathcal{A}$, the sets $\mathcal{B}[2A]$ and $\mathcal{B}[2A^{\prime}]$ are disjoint, where $2A$ denotes the 2-fold dilate of $A$. In practice, $\mathcal{A}$ and $\mathcal{B}$ will be collections of $\rho$-tubes and $\delta$-tubes respectively, for some $\rho>\delta$.
@@
Let $C\geq 1,\delta>0$. We say that a set $\mathbb{T}$ of $\delta$-tubes in $\mathbb{R}^{n}$ satisfies the Convex Wolff Axioms at every scale with error $C$ if the tubes in $\mathbb{T}$ are essentially distinct, and for every $\rho_{0}\in[\delta,1]$, there exists $\rho\in[\rho_{0},C\rho_{0})$ and a set of $\rho$-tubes $\mathbb{T}_{\rho}$ that satisfies the following properties:
@@
    For each $T_{\rho}\in\mathbb{T}_{\rho}$, the unit rescaling of $\mathbb{T}[T_{\rho}]$ relative to $T_{\rho}$ satisfies the Convex Wolff Axioms with error $C$.
As the name suggests, Definition 2.12 is a multi-scale property. In particular, if $\mathbb{T}$ satisfies the Convex Wolff Axioms at every scale, then we can cover $\mathbb{T}$ by fatter $\rho$-tubes; both the $\rho$-tubes and the (rescaled) tubes from $\mathbb{T}$ inside each $\rho$-tube will again satisfy the Convex Wolff Axioms at every scale. The precise statement is as follows.